OpenStudy (anonymous):

Graph the inequality -2x - y - 1

OpenStudy (jdoe0001):

inequality?

OpenStudy (anonymous):

yes thats what it says

OpenStudy (jdoe0001):

this is what I see => \(\huge -2x - y - 1 \)

OpenStudy (anonymous):

Yes i can even take a picture of my wkst...

OpenStudy (jdoe0001):

hmmm. ok... I take a picture then

OpenStudy (jdoe0001):

because is missing an inequality sign

OpenStudy (anonymous):

ok wait like few min

OpenStudy (anonymous):

it's #7

OpenStudy (jdoe0001):

hmm, is a typo then, if I were to make anything out of that, I'd think is \(\huge -2x\ge -y-1\ or \ -2x \le -y-1\)

OpenStudy (anonymous):

alright then Ill skip it.. can you help me with #6 then

OpenStudy (jdoe0001):

well, you'd solve for "y" first what would that give you?

OpenStudy (anonymous):

so divide by -2?

OpenStudy (jdoe0001):

yes, keep in mind that, when multiplying/dividing/exponentiazling by a negative value, you need to \(\bf \text{flip}\) the sign

OpenStudy (anonymous):

\[y \ge x/2 + 1\]

OpenStudy (jdoe0001):

yeap, so in essence the graph will be, using the line of y = x/2 + 1 the line drawn will be SOLID, because is using \(\bf \ge\) for > or < is a DASHED line

OpenStudy (jdoe0001):

is a straight line, so all you need is 2 points

OpenStudy (anonymous):

how do i graph it? the x/2 part

OpenStudy (jdoe0001):

well, one of the easiest way to get 2 points, is set x = 0, then set y = 0 so \(\bf y \ge \cfrac{x}{2}+1\\ \quad \\ \quad \\ x =0,\qquad y = \cfrac{0}{2}+1\implies y =1\quad (0, 1)\\ \quad \\ y=0, \qquad 0 = \cfrac{x}{2}+1 \implies -2 = x\quad (-2, 0)\)

OpenStudy (anonymous):



OpenStudy (anonymous):

do i need to find the points or can i just look at the equation?

OpenStudy (jdoe0001):

yeap, that's the line so for the shading, we do some testing on a point NOT IN THE LINE say let's use (0, 0) is not in the line, so we'll use that \(\bf y \ge \cfrac{x}{2}+1\\ \quad \\ \quad \\ (0,0)\\ \quad \\ y \ge \cfrac{x}{2}+1\implies 0 \ge \cfrac{0}{2}+1\implies 0\ge 1\) so, is 0 really greater or equal to 1? well, not

OpenStudy (jdoe0001):

so the region where (0,0) is at, is the FALSE area so the TRUE area will be on the other side of the line, and that's the one that gets the shading

OpenStudy (anonymous):

do i need to find the points or can i just look at the equation?

OpenStudy (jdoe0001):

for the shading?

OpenStudy (anonymous):

no to graph it

OpenStudy (anonymous):

is x/2 the same as 2x?

OpenStudy (jdoe0001):

hmm, no.... x/2 is one half, 2x is twice as much

OpenStudy (jdoe0001):

to graph it? you just did :)

OpenStudy (jdoe0001):

the graph is just that one line you drew

OpenStudy (anonymous):

then why did you find the (0,1) and (-2,0)

OpenStudy (jdoe0001):

ohh that's for the shading to show in the graph, where the inequality applies that is, what part of the graph the values for which that inequality http://www.regentsprep.org/regents/math/algebra/ae85/grineqa.htm

OpenStudy (jdoe0001):

hhhmmm shoot I misunderstood.. anyhow... the ((0,1) and (-2,0) part was just to get 2 points to make the line, but you did just fine

OpenStudy (jdoe0001):

the (0,0) testing part is for the shading of it

OpenStudy (anonymous):

ohh so my graph is okay too?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

alright thanks!! would you mind helping me with one more thing?

OpenStudy (anonymous):

@JuanitaM

OpenStudy (anonymous):

@jdoe0001